A population of birds has a mean beak-size of 5cm. A drought occurs such that a...

  1. A population of birds has a mean beak-size of 5cm. A drought occurs such that a large beaked birds survived better than small-beaked birds. The large-beaked birds that survived had an average beak size of 7 cm, and their offspring had an average beak size of 6cm. What is the selection differential?
    1. 1
    2. 2
    3. 3
    4. 4
    5. None of the abov

* Selection differential= large beaked birds that survive – population mean beak sized

  1. What is the response to selection in the above question?
    1. 1
    2. 2
    3. 3
    4. 4
    5. None of the above

* R= H2 X S

R- change in the mean over one complete generation

S= is equal to the difference in mean trait values between the selected individual and the entire population

  1. Calculate heritability from question 18 above.
    1. 0.25
    2. 0.33
    3. 0.50
    4. 0.75
    5. None of the abov

*** H2= S/R

Could you help me, know how to solve these questions???
Answer #1

In the first part of question:

Selection differential= large beaked birds that survive – population mean beak sized

                                  =7-5=2

The Grants observed evolution happening: the average beak size increased after a drought, only large-beaked birds were able to crush the bigger seeds and survive to make the next generation, when wet periods returned smaller beaks prevailed.

in the second part:

response to selection=large beaked birds that survive – offspring had an average beak size

                                  =7-6=1

heritability=1/2=.5